# MEMERAG: A Multilingual End-to-End Meta-Evaluation Benchmark for Retrieval Augmented Generation

### v1

@@ Abstract
In this work, we develop a Multilingual End-to-end Meta-Evaluation RAG benchmark (MEMERAG). Our benchmark builds on the popular MIRACL dataset, using native-language questions and generating responses with diverse large language models (LLMs), which are then assessed by expert annotators for faithfulness and relevance. We describe our annotation process and show that it achieves high inter-annotator agreement. We then analyse the performance of the answer-generating LLMs across languages as per the human evaluators. Finally we apply the dataset to our main use-case which is to benchmark multilingual automatic evaluators (LLM-as-a-judge). We show that our benchmark can reliably identify improvements offered by advanced prompting techniques and LLMs  [^1] We release https://anonymous our benchmark to support the community developing accurate evaluation methods for multilingual RAG systems..
@@ 3.1 Question Selection
The questions in the MIRACL dataset were generated by humans based on prompts. This leads to questions that may be answerable by the prompt but may have ambiguity outside of that context. In particular, we identified as problematic the questions for which the right answer can change over time. For example “Who is the president of Spain?” or “How old is Drake Hogestyn?”. In a RAG setting, different passages may have been written at different times and provide conflicting context. Additionally, the time of reference is usually not explicit in the question. To remove those complications we automatically filtered out time-dependent questions across all languages444 See Appendix F for all prompts used during dataset construction. . We combined the train and dev splits of the MIRACL dataset, which corresponds to a total of 3,662, 305, 2,810, 1,486 and 1,519 questions respectively for EN, DE, ES, FR and HI. Of those, 244, 13, 120, 64 and 86 were identified as time-dependent and filtered out.
@@ 3.5 Annotation Process
Table 2 summarizes the IAA per language for faithfulness and relevance labels assigned by 3 annotators. We report IAA using Gwet’s AC1 Gwet (2008) and Fleiss Kappa Fleiss (1971). We observe high agreement for faithfulness (0.84-0.93 Gwet’s AC1 and 0.70-0.88 Fleiss Kappa) and even higher agreement for relevance (0.95-1.0 Gwet’s AC1 and 0.63-1.0 Fleiss Kappa). This shows that the annotators are aligned and indicates a high quality of the annotations. Comparing to previous work, Tang et al. (2024b) report a Fleiss Kappa of 0.34-0.42 on faithfulness labels which they deem fair to moderate agreement. Note that a direct comparison with this work is not possible as they deal with different tasks, nevertheless the high IAA we are reporting is a testament of the effectiveness of our flow chart-based annotation design. Further details on IAA including fine-grained explanatory labels can be found in Appendix A, Table 6.
@@ 5.1 Experimental Setup
Our benchmark is composed of multiple tasks defined by the annotation dimension and subset considered. On the annotation dimension, we focus our experiments on the coarse-grained faithfulness dimension. This dimension is more challenging than relevance as highlighted by the lower IAA, while retaining a high-enough IAA to make for a trustworthy benchmark. We invite benchmark users to also experiment on the other dimensions provided by the dataset depending on their use case. Note that we remove the sentences labelled as Challenging to determine by human annotators. On the annotation subset, we first distinguish the multilingual task which uses the full dataset and the monolingual task, which only considers a single language. Those task can then be further broken down at the fine-grained level by considering the subset of sentences with a certain fine-grained label. Performance is evaluated with Balanced Accuracy (BAcc) (see Appendix  E for definition), with equal weights on each coarse-grained label and language. We conduct significance testing using permutation tests Good (2013), with further details in Appendix D.
@@
To demonstrate how the benchmark can be used to select the appropriate prompt, we experiment with multiple prompting strategies from simple to advanced techniques, starting with zero-shot prompting, where LLMs directly classify statements as Supported or Not supported. We then implement chain-of-thought (COT) prompting Wei et al. (2022), which incorporates an intermediate reasoning step. While these basic prompting strategies provide a good starting point for evaluation, our initial experiments revealed limitations in their ability to capture the nuanced requirements of our specific task. Without explicit guidelines in the prompt, automatic evaluators rely on their “world knowledge”, which may not align with the specific requirements of the evaluation task. To overcome this, we add instructions from the annotation guidelines (AG) in the prompt. Adding annotation guidelines provides clear criteria for what constitutes Supported versus Not Supported sentences, reducing ambiguity in the evaluation process. The various prompts are presented in Appendix G.
@@
Table 5: Reference baselines for the multilingual task on coarse-grained faithfulness. Balanced accuracy (BAcc) of the automatic evaluators averaged across the 5 languages (EN, DE, ES, FR, HI) using zero-shot (ZS), chain-of-thought (COT), annotation guidelines (AG) and AG+COT prompting strategies. Bold indicates best performance for the column, $\dagger$ indicates results not statistically different from the best (p > 0.05). Additional results on monolingual tasks and standard errors can be found in Appendix D, Tables 8-10.
@@ 5.2 Experimental Results
Figure  2, shows the performance per language of various automatic evaluators with a fixed prompt (AG + COT), which allows us to select the best model for each language. We observe that GPT-4o mini performs best in English. For the rest of the languages Qwen 2.5 32B performs the best however, the results are not statistically different from GPT-4o mini. Our benchmark also provides users with the capability to conduct detailed, fine-grained analyses of model performance across various dimensions of faithfulness. The breakdown of automatic evaluation performance by error type is shown in Appendix  I.
@@ Appendix A Human Annotation Guidance
Upon careful analysis of the annotator disagreements in the pilot and further rounds of calibration, we developed the annotation workflow shown in Figure 3. The key improvements were: (i) add a Challenging to determine label, (ii) have two levels of labels, a coarse-grained level (Supported, Not supported, Challenging to determine) and a fine-grained level for precision on the mistakes, (iii) enforce annotators to follow a specific reasoning with a flow chart, (iv) use numbers for the fine-grained level rather than labels which could be misinterpreted. Those guidelines allowed to reach significantly higher IAA on faithfulness (Gwet’s AC1 0.81 coarse-grained). Likewise we iterated on the relevance labels, starting from Must have, Nice to have, or Irrelevant and converging to the more explicit Directly answers the question, Adds context to the answer, or Unrelated to the question. This also increased the IAA significantly. The screenshot of the user interface used by human annotators is shown in Figure  4
@@
Table 6: Inter-annotator agreement (IAA) on the faithfulness and relevance dimensions with 3 annotators, for coarse-grained and fine-grained levels. Annotations are at the answer sentence level (#S number of answer sentences is provided).
Appendix B Detailed numbers on MEMEREG dataset
@@
Appendix C Fine grained automatic annotation analysis
@@
Appendix D Statistical Significance Test Details
@@
Appendix E Balanced Accuracy
@@
Appendix F Prompts used for dataset construction
@@
F.1 Time-dependent answer filtering
@@
F.2 Highlighting relevant segments
@@
F.3 Answer generation prompt
@@
Appendix G Prompts for automated evaluation
@@
Appendix H More Results on Meta-Evaluators
@@
Appendix I Fine grained Analysis
@@
To demonstrate the benchmark’s capability for fine-grained analysis, Table 11 breaks down automatic evaluation performance by error type for two models: GPT-4o mini and Llama 3.2 90B, both evaluating French answers using AG + COT prompt. The upper section of the table shows the distribution of errors when the ground truth class is Supported, categorized into logical conclusion, direct paraphrase, or other correct categories. The lower section details the types of mistakes made by each model for the Not supported category. We observe that both models show a similar pattern for the Supported category, with logical conclusions being the most common (69.8% for GPT-4o mini and 76.0% for Llama 3.2 90B), followed by direct paraphrases. For the Not supported category, both models struggle most with detecting “adding new information” (32.1% and 40.0% of mistakes, respectively) and “nuance shifts” (27.2% and 18.5%). The ranking of error types is consistent across both models, despite their different overall BAcc (73.7% for GPT-4o mini vs 63.2% for Llama 3.2 90B see Table  10 in Appendix), suggesting similar challenges in evaluation. The benchmark’s fine-grained labeling system provides a detailed view of evaluation challenges across languages. By categorizing different types of faithfulness violations, it reveals which specific errors are harder to detect for each model, offering insights into both the strengths and limitations of automated evaluation methods.

### v3

@@ Abstract
In this work, we develop a Multilingual End-to-end Meta-Evaluation RAG benchmark (MEMERAG). Our benchmark builds on the popular MIRACL dataset, using native-language questions and generating responses with diverse large language models (LLMs), which are then assessed by expert annotators for faithfulness and relevance. We describe our annotation process and show that it achieves high inter-annotator agreement. We then analyse the performance of the answer-generating LLMs across languages as per the human evaluators. Finally we apply the dataset to our main use-case which is to benchmark multilingual automatic evaluators (LLM-as-a-judge). We show that our benchmark can reliably identify improvements offered by advanced prompting techniques and LLMs  [^1] We release https://github.com/amazon-science/MEMERAG our benchmark to support the community developing accurate evaluation methods for multilingual RAG systems..
@@ 3.1 Question Selection
The questions in the MIRACL dataset were generated by humans based on prompts. This leads to questions that may be answerable by the prompt but may have ambiguity outside of that context. In particular, we identified as problematic the questions for which the right answer can change over time. For example “Who is the president of Spain?” or “How old is Drake Hogestyn?”. In a RAG setting, different passages may have been written at different times and provide conflicting context. Additionally, the time of reference is usually not explicit in the question. To remove those complications we automatically filtered out time-dependent questions across all languages444 See Appendix G for all prompts used during dataset construction. . We combined the train and dev splits of the MIRACL dataset, which corresponds to a total of 3,662, 305, 2,810, 1,486 and 1,519 questions respectively for EN, DE, ES, FR and HI. Of those, 244, 13, 120, 64 and 86 were identified as time-dependent and filtered out.
@@ 3.5 Annotation Process
Table 2 summarizes the IAA per language for faithfulness and relevance labels assigned by 3 annotators. We report IAA using Gwet’s AC1 Gwet (2008) and Fleiss Kappa Fleiss (1971). We observe high agreement for faithfulness (0.84-0.93 Gwet’s AC1 and 0.70-0.88 Fleiss Kappa) and even higher agreement for relevance (0.95-1.0 Gwet’s AC1 and 0.63-1.0 Fleiss Kappa). This shows that the annotators are aligned and indicates a high quality of the annotations. Comparing to previous work, Tang et al. (2024b) report a Fleiss Kappa of 0.34-0.42 on faithfulness labels which they deem fair to moderate agreement. Note that a direct comparison with this work is not possible as they deal with different tasks, nevertheless the high IAA we are reporting is a testament of the effectiveness of our flow chart-based annotation design. Further details on IAA including fine-grained explanatory labels can be found in Appendix B.
@@ 5.1 Experimental Setup
Our benchmark is composed of multiple tasks defined by the annotation dimension and subset considered. On the annotation dimension, we focus our experiments on the coarse-grained faithfulness dimension. This dimension is more challenging than relevance as highlighted by the lower IAA, while retaining a high-enough IAA to make for a trustworthy benchmark. We invite benchmark users to also experiment on the other dimensions provided by the dataset depending on their use case. Note that we remove the sentences labelled as Challenging to determine by human annotators. On the annotation subset, we first distinguish the multilingual task which uses the full dataset and the monolingual task, which only considers a single language. Those task can then be further broken down at the fine-grained level by considering the subset of sentences with a certain fine-grained label. Performance is evaluated with Balanced Accuracy (BAcc) (see Appendix F for definition), with equal weights on each coarse-grained label and language. We conduct significance testing using permutation tests Good (2013), with further details in Appendix E.
@@
To demonstrate how the benchmark can be used to select the appropriate prompt, we experiment with multiple prompting strategies from simple to advanced techniques, starting with zero-shot prompting, where LLMs directly classify statements as Supported or Not supported. We then implement chain-of-thought (COT) prompting Wei et al. (2022), which incorporates an intermediate reasoning step. While these basic prompting strategies provide a good starting point for evaluation, our initial experiments revealed limitations in their ability to capture the nuanced requirements of our specific task. Without explicit guidelines in the prompt, automatic evaluators rely on their “world knowledge”, which may not align with the specific requirements of the evaluation task. To overcome this, we add instructions from the annotation guidelines (AG) in the prompt. Adding annotation guidelines provides clear criteria for what constitutes Supported versus Not Supported sentences, reducing ambiguity in the evaluation process. The various prompts are presented in Appendix H.
@@
Table 5: Reference baselines for the multilingual task on coarse-grained faithfulness. Balanced accuracy (BAcc) of the automatic evaluators averaged across the 5 languages (EN, DE, ES, FR, HI) using zero-shot (ZS), chain-of-thought (COT), annotation guidelines (AG) and AG+COT prompting strategies. Bold indicates best performance for the column, $\dagger$ indicates results not statistically different from the best (p > 0.05). Additional results on monolingual tasks and standard errors can be found in Appendix E, Tables 10-12.
@@ 5.2 Experimental Results
Figure  2, shows the performance per language of various automatic evaluators with a fixed prompt (AG + COT), which allows us to select the best model for each language. We observe that GPT-4o mini performs best in English. For the rest of the languages Qwen 2.5 32B performs the best however, the results are not statistically different from GPT-4o mini. Our benchmark also provides users with the capability to conduct detailed, fine-grained analyses of model performance across various dimensions of faithfulness. The breakdown of automatic evaluation performance by error type is shown in Appendix J.
@@ Appendix A Human Annotation Guidance
Upon careful analysis of the annotator disagreements in the pilot and further rounds of calibration, we developed the annotation workflow shown in Figure 3. The key improvements were: (i) add a Challenging to determine label, (ii) have two levels of labels, a coarse-grained level (Supported, Not supported, Challenging to determine) and a fine-grained level for precision on the mistakes, (iii) enforce annotators to follow a specific reasoning with a flow chart, (iv) use numbers for the fine-grained level rather than labels which could be misinterpreted. Those guidelines allowed to reach significantly higher IAA on faithfulness (Gwet’s AC1 0.81 coarse-grained). Likewise we iterated on the relevance labels, starting from Must have, Nice to have, or Irrelevant and converging to the more explicit Directly answers the question, Adds context to the answer, or Unrelated to the question. This also increased the IAA significantly. The screenshot of the user interface used by human annotators is shown in Figure  4.
Appendix B IAA and Extended Dataset
MEMERAG
@@ Appendix B IAA and Extended Dataset
Table 6: Inter-annotator agreement (IAA) on the faithfulness and relevance dimensions with 3 annotators (MEMERAG subset), for coarse-grained and fine-grained levels. Annotations are at the answer sentence level (#S number of answer sentences is provided).
MEMERAG-Ext
Lang #S IAA (Gwet’s AC1)
Faithfulness Relevance
Coarse Fine Coarse Fine
EN 226 0.83 0.41 1.00 0.9
DE 272 0.75 0.47 0.92 0.79
ES 276 0.91 0.39 1.00 0.97
FR 370 0.72 0.53 0.99 0.8
HI 208 0.91 0.33 0.98 0.92
Table 7: Inter-annotator agreement (IAA) on the faithfulness and relevance dimensions with 5 annotators (MEMERAG-Ext), for coarse-grained and fine-grained levels. Annotations are at the answer sentence level (#S number of answer sentences is provided).
As explained in Section 3.5, the IAA on the MEMERAG dataset was measured by triple annotation on a random subset of 10 questions per language. Detailed IAA metrics, at both coarse- and fine-grained levels, can be found in Table 6. We observe that IAA is high for faithfulness at coarse-grained level, and for relevance at both levels. This prompted us to 1) focus the experiments described in this paper to those dimensions with high IAA, and 2) investigate the low IAA for fine-grained faithfulness.
The disagreements highlighted by low IAA are due to ambiguity in the annotation guidelines. The fact that such ambiguity remained after significant work on refining the guidelines and guiding annotators with UI elements hints at the irreducible subjectivity of the task. Subjectivity manifests itself as a non-trivial probability distribution of the annotation label obtained by drawing a random human annotator for a given question. It would be of interest in the development of automated annotators to compare this distribution to that of the annotations of a stochastic LLM-based annotator.
To investigate those aspects, we created an extended dataset, called MEMERAG-Ext. The dataset is composed of 150 question-context pairs per language, extracted similarly to MEMERAG with the process described in Section 3. The questions are disjoint between the two datasets, except for German where there is some overlap due to the limited number of questions in MIRACL. The annotation process then differed in two ways: 1) annotations were obtained for all question from 5 expert annotators, and 2) The Challenging to determine label was not an option for annotators to force meaningful labels even at the cost of disagreements. IAA metrics for the MEMERAG-Ext dataset can be found in Table 7. Those results are in line with the IAA observed in MEMERAG, which confirms its quality. Detailed investigation of the disagreements between human annotators, and comparison to automated annotators are left as future work. We release this extendended dataset along with the main dataset to promote related work by the research community.
Lang #Q Answer Context
#S Avg. #W Avg. #W
EN 150 226 27.5 591.4
DE 150 272 30.0 456.8
ES 150 276 39.1 490.8
FR 150 370 48.1 453.2
HI 150 208 23.0 576.6
Total 750 1352
Table 8: General statistics of the MEMERAG-Ext dataset. #Q: number of questions, #S: number of sentences, #W: number of words. Each answer is annotated at the sentence level, leading to 1,352 total sentences annotated by experts for faithfulness and relevance.
Appendix C Detailed numbers on the MEMERAG dataset
@@
Appendix D Fine grained automatic annotation analysis
@@
Appendix E Statistical Significance Test Details
@@
Appendix F Balanced Accuracy
@@
Appendix G Prompts used for dataset construction
@@
G.1 Time-dependent answer filtering
@@
G.2 Highlighting relevant segments
@@
G.3 Answer generation prompt
@@
Appendix H Prompts for automated evaluation
@@
Appendix I More Results on Meta-Evaluators
@@
Appendix J Fine grained Analysis
@@
To demonstrate the benchmark’s capability for fine-grained analysis, Table 13 breaks down automatic evaluation performance by error type for two models: GPT-4o mini and Llama 3.2 90B, both evaluating French answers using AG + COT prompt. The upper section of the table shows the distribution of errors when the ground truth class is Supported, categorized into logical conclusion, direct paraphrase, or other correct categories. The lower section details the types of mistakes made by each model for the Not supported category. We observe that both models show a similar pattern for the Supported category, with logical conclusions being the most common (69.8% for GPT-4o mini and 76.0% for Llama 3.2 90B), followed by direct paraphrases. For the Not supported category, both models struggle most with detecting “adding new information” (32.1% and 40.0% of mistakes, respectively) and “nuance shifts” (27.2% and 18.5%). The ranking of error types is consistent across both models, despite their different overall BAcc (73.7% for GPT-4o mini vs 63.2% for Llama 3.2 90B see Table  12 in Appendix), suggesting similar challenges in evaluation. The benchmark’s fine-grained labeling system provides a detailed view of evaluation challenges across languages. By categorizing different types of faithfulness violations, it reveals which specific errors are harder to detect for each model, offering insights into both the strengths and limitations of automated evaluation methods.